sirp:

A submarine cruises underwater at 10km/h and on the surface at 16km/h. the submarine travels a distance of 160km in the 12.5h. how long did it travel underwater? how long did it travel on the surface?

Mercury:

you can write a system of equations, one for the total time and one for the total distance let u = # of hours spent underwater and s = # of hours spent on the surface it says that the total time travelled was 12.5 hours, so s + u = 12.5 now, for distance --> distance = speed*time therefore, underwater distance = (underwater speed)(time spent underwater) same logic, surface distance = (surface speed)(time spent on surface) which translates to 10u + 16s = 160

Mercury:

so, now your system is: u + s = 12.5 and 10u + 16s = 160 solve for s